OpenStudy (anonymous):

Find all critical numbers and use the Second Derivative Test to determine all local extrema. y = x^2 - (16)/(x)

OpenStudy (anonymous):

@satellite73 can u help?

OpenStudy (anonymous):

You must mean first derivative test, unless you're also checking for intervals of concavity? Do you know the derivative of the given function? \[y=x^2-\frac{16}{x}\]

OpenStudy (anonymous):

yea i found he derivatuve its \[2x + \frac{ 16 }{ x^2 } \]

OpenStudy (anonymous):

Okay, you know how to find the critical points? Setting \(y'=0,\) etc?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got x = -2

OpenStudy (anonymous):

Good. But keep in mind that \(y'\) is undefined when \(x=0\), so that would be another critical point.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So when applying the derivative test, you have the intervals \[(-\infty,-2), (-2,0),\text{ and }(0,\infty).\] Find the sign of the derivative on each of these intervals.

OpenStudy (anonymous):

okay how do i do that?

OpenStudy (anonymous):

Take any value of x that lies on one of the intervals. For example, take \(x = -1\). This lies in the second interval \((-2,0)\). \[y'(1)=2(-1)+\frac{16}{(-1)^2}=-18<0\] Since the derivative is negative on this interval, you know that \(y\) is decreasing on this interval.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so so -infi, -2 i should take -1 right?

OpenStudy (anonymous):

No, \(-1\) isn't in that interval. You have to pick a number less than \(-2\).

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so -3?

OpenStudy (anonymous):

\[\large y = x^2 -\frac{16}{x}\] \[\large \frac{dy}{dx}=2x+\frac{16}{x^2}\] \[\large \frac{d^2y}{dx^2}=2-\frac{16}{x^3}\] At S.P's dy/dx=0 \[\large \frac{2x^3+16}{x^2}=0\] \[\large 2x^3=-16\] \[\large x^3=-8\] \[\large x=-2\] Sub the x-value into the original equation to find the y-coordinate. \[\large y=(-2)^2-\frac{16}{-2}\] \[\large y=12\] Sub the x-value into the second derivative to determine whether it is a maximum or minimum turning point at that particular point. \[\large y">0\] \[\large =>Minimum T.P (Turning Point) at (-2, 12)\] At I.P's y"=0 \[\large \frac{2x^3-16}{x^3}=0\] \[\large x^3=8\] \[\large x=2\] Draw a table to see whether x=2 is an inflexion point.